OpenStudy (anonymous):

find the equation of the line that passes through the point(4,6) and the x intercept of 3x-2y +9=0 by the way i need explanation and working out thank you!

OpenStudy (anonymous):

you need the find the x-intercept of \[3x-2y+9=0\] and then you will have two points with which to find the equation of the line

OpenStudy (anonymous):

to find it set y = 0 and solve \[3x+9=0\] \[3x=-9\] \[x=-3\] so the x-intercept is (-3,0)

OpenStudy (anonymous):

now you have two points (-3,0) and (4,6) the slope is \[\frac{6-0}{4-(-3)}=\frac{6}{7}\]

OpenStudy (anonymous):

use the point - slope formula to find the equation of the line: \[y-y_1=m(x-x_1)\] \[y-0=\frac{6}{7}(x-(-3))\] \[y = \frac{6}{7}x+\frac{18}{7}\]

OpenStudy (anonymous):

thank you! the answer provided was 7y=6x+18 which is the same it all makes sense now! ^^